# Visual Agents as Fast and Slow Thinkers

### v3

@@ 2 Methods
Notation. The integration of components in visual agents $\mathcal{F}$ (based on the Large Language Model) typically involves a visual encoder, denoted as $\mathcal{E}_{V}$, a nature language encoder, represented by $\mathcal{E}_{L}$, and a Language Language Model such as Vicuna (Chiang et al., 2023). Initially, the visual agent is presented with an image $\mathcal{I}$ and an accompanying textual prompt $\mathcal{Q}$, which could be a question or instruction. Then the visual agent combines these multimodal tokens into a united space. Finally, the visual agent outputs a textual response $R$ given the textual and image input. The generation process can be expressed as Eq. 1:
@@ 2.1 FaST
System Switch. Current works on visual agents mostly rely on visual question-answering data, which gives direct answers (System 1) after inquiry as Eq. 1. However, attempting to answer questions directly in this way can compromise the reliability of the responses. Agents tend to hallucinate over questions that require more deliberate reasoning and visual details. To reduce hallucination and make the model reliable, we utilize a system switch trigger to tell when to require more visual information. Specifically, for a question $\mathcal{Q}$ and an image $\mathcal{I}$, we define a MLLM with switch adapter $\mathcal{S}$ and formulate the fast $\mathcal{F}_{fast}$ and slow thinking process $\mathcal{F}_{slow}$. When the query is easy, the frame does not need the switch adapter $\mathcal{S}_{adapter}$ and only output result $\mathcal{R}$ by $\mathcal{F}_{fast}$ as Eq. 2.
@@ 2.2 Preliminary
Hierarchical Reasoning. We use a top-down scheme to reason over multi-scale granularity images effectively in order to reason and take advantage of world knowledge progressively. Similar to humans would look for some context clue to find specific objects relating to questions and zoom in if they think the answer lies in a particular region, we model this process with system switch adapters as Eq. 2.2 to focus on the context clue $\mathcal{C}_{clue}$ generated from the switch adapter as Eq. 3.
We denote the MLLM as a proposal adapter $\mathcal{P}_{adapter}$ (visual agent). In System 2, FaST uses many visual agents to accomplish hierarchical reasoning. The frame tries to narrow down the search space by using the question $\mathcal{Q}$ and the previously obtained clue $\mathcal{C}_{clue}$ to let the proposal adapter output a region $Region$ that aligns with the question and the context clue as Eq. 4.
@@
After getting the region, the visual agents $\mathcal{P}_{adapter}$ will be asked to focus on a more specific target with a bounding box $[Bboxes]$ complemented by the context clue $\mathcal{C}_{clue}$ and region $Region$ get from Eq. 4. This process can reveal the step-by-step reasoning and be modeled as Eq. 5.
@@
When we have a more specific target proposal(bounding box $[Bboxes]$), FaST will apply a fine-grained pixel-level mask decoder $\mathcal{P}_{Seg}$ as Eq. 6 to output the specific mask part $[Mask]$ of the target proposal $[Bboxes]$ to focus on as Eq. 6. We name this whole process from $Region$ to $[Mask]$ chain of evidence as Remark. 2.5 similar to thinking more and more deeply by humans.
@@
After getting the target proposal (bounding box $[Bboxes]$) from context clue $\mathcal{C}_{clue}$ with proposal adapter and specific mask part $[Mask]$ by missing objects with seg adapter, a chain of evidence is constructed as Remark 2.5 and Fig. 3. Our FaST framework then summarizes all this information ($\mathcal{I}$ and $\mathcal{Q}$) and the chain of evidence with switch adapter to give the final correct reasoning answer $Ans$ as Eq. 7
@@ 2.3 Implementation Details
$\bullet$ Chain of Evidence. When we apply the chain of evidence as Remark 2.5 in the LLM to get the answer as the final step like Eq. 7. The whole sequence of the chain of evidence is too long to load in the $\mathcal{F}_{Slow}$. So FaST needs a visual sampler based on cross-attention that is trained to decrease the number of image tokens to a suitable length (from 256 to 32), apart from MLP projection.
@@ 4 Related works
LLM as Visual Agents. With the capabilities that LLMs have demonstrated in language understanding and generation (Ouyang et al., 2022; OpenAI, 2024; Zheng et al., 2023; Touvron et al., 2023a; b; Wang et al., 2024a; Hua et al., 2024; Mei et al., 2024; Jin et al., 2024a), the research community has progressed to explore how LLMs can be enhanced with vision input for multimodal tasks as visual agents (Alayrac et al., 2022; Driess et al., 2023; Li et al., 2023a; Ge et al., 2023; Dai et al., 2024; Liu et al., 2023b; 2024a; Lin et al., 2024). There are two paradigms for LLM-based visual agents: end-to-end based and tool-using visual agents. Following the principle of instruction tuning, end-to-end visual agents are trained with a curated visual instruction tuning dataset to digest features from multi-modality, unlocking the capability to answer visual questions (Huang et al., 2023; Luo et al., 2023; Zhu et al., 2023a; Bai et al., 2023; Zhang et al., 2023b; c; Chen et al., 2024a; Ye et al., 2023; Singh et al., 2019). For other visual tasks ($e.g.$, Segmentation, Detection, etc), end-to-end trained tailored agents can further perform downstream tasks (Pi et al., 2023; Peng et al., 2024; Lai et al., 2024; Chen et al., 2023b; Wang et al., 2023c; Dai et al., 2024; Wang et al., 2024b; 2023b; Jiang et al., 2023; Chen et al., 2023a). Recent research has focused on leveraging improved vision encoders and fostering more detailed visual understanding, yielding promising results (Fan et al., 2024; Xu et al., 2024a; Shi et al., 2024). While these approaches can be implemented with direct instruction tuning data, they represent a ‘System 1’ type of training. This type of training primarily relies on the dataset’s quality and tends to provide direct answers that are prone to hallucinations, a consequence inherent to the nature of System 1 instruction tuning data. For the second paradigm, tool-using models are built on top of a frozen LLM with access to pretrained visual perception tools (Surís et al., 2023; Shen et al., 2023; Lu et al., 2023a). In this scenario, the LLM first selects visual tools and then decides by thoroughly analyzing the fine-grained information extracted by visual tools (Lu et al., 2023a; You et al., 2023; Wu et al., 2024). While external visual tools enhance the interpretability of the reasoning process, their complexity can introduce inaccuracies. Moreover, the abundance of information generated during reasoning may overshadow key details relevant to the query, resulting in incorrect answers.
@@
System 2 in AI. Recently, LLMs have been engineered to produce text that mimics the step-by-step reasoning process characteristic of human cognition, akin to the analytical and deliberate thought processes associated with what is termed as System 2 in the human cognition process (Qiao et al., 2023; Huang & Chang, 2023a; Wang et al., 2023a; Shaikh et al., 2023; Shao et al., 2024). The systematic approach to problem-solving is a hallmark across various domains, including mathematical word problems (Kojima et al., 2022; Wang et al., 2023d; Lightman et al., 2023; Cobbe et al., 2021; Liu et al., 2023c; Zhu et al., 2023b; Lu et al., 2023b), logical reasoning (Yao et al., 2023d; a; Besta et al., 2024; Wen et al., 2023; Lei et al., 2023; Cheng et al., 2024; Jin et al., 2024b), and multi-modal reasoning (Chen et al., 2024d; You et al., 2023; Wu & Xie, 2024). In Explainable AI, this systematic method is emulated by the model as it generates a text-based elucidation of its reasoning and decision-making process through step by step reasoning process ($e.g.$, chain of thought) (Han et al., 2024; Zhao et al., 2024; Jacovi & Goldberg, 2020; Hua & Zhang, 2022). However, it is crucial to recognize that LLMs, while powerful, are not exempt from encountering challenges when facing complex problems. One such challenge is the issue of hallucination (Zhou et al., 2024; Cui et al., 2023; Li et al., 2023b; Zhang et al., 2023a; Chen et al., 2024c; Guan et al., 2024a), which can distort the model’s reasoning process and lead to inaccuracies in the explanations provided. Initially, LLM reasoning is seen as only a linear chain of thoughts, where each step in the reasoning process is clearly articulated. As models evolve, they adopt more complex structures like hierarchical trees (Geng et al., 2023; Yao et al., 2023b) and intricate graphs (Besta et al., 2024), which enable them to handle much more complex problems but also restrict their general applicability because of increased topological complexity (Yao et al., 2023b; a; Besta et al., 2024; Lei et al., 2023; Wen et al., 2023). Moreover, these complex structures can lead to errors propagating through the model’s reasoning, causing a cascade of mistakes (Xu et al., 2024b). To counter this, incorporating feedback from intermediate reasoning steps and employing iterative refinement, which is similar to human reflection, could help mitigate errors (Chu et al., 2023; Tong et al., 2024b; Guan et al., 2024b; Madaan et al., 2023; Yuan et al., 2024; Wu & Xie, 2024). In unsupervised scenarios, such feedback is vital for enhancing the reasoning capabilities of LLMs and reducing errors (Yao et al., 2023c).

### v4

@@ 2 Methods
Notation. The integration of components in visual agents $\mathcal{F}$ (based on the Large Language Model) typically involves a visual encoder, denoted as $\mathcal{E}_{V}$, a nature language encoder, represented by $\mathcal{E}_{L}$, and a Language Language Model such as Vicuna (Chiang et al., 2023). Initially, the visual agent is presented with an image $\mathcal{I}$ and an accompanying textual prompt $\mathcal{Q}$, which could be a question or instruction. Then the visual agent combines these multimodal tokens into a united space. Finally, the visual agent outputs a textual response $R$ given the textual and image input. The generation process can be expressed as Equation 1:
@@ 2.1 FaST
System Switch. Current works on visual agents mostly rely on visual question-answering data, which gives direct answers (System 1) after inquiry as Eq. 1. However, attempting to answer questions directly in this way can compromise the reliability of the responses. Agents tend to hallucinate over questions that require more deliberate reasoning and visual details. To reduce hallucination and make the model reliable, we utilize a system switch trigger to tell when to require more visual information. Specifically, for a question $\mathcal{Q}$ and an image $\mathcal{I}$, we define a MLLM with switch adapter $\mathcal{S}$ and formulate the fast $\mathcal{F}_{fast}$ and slow thinking process $\mathcal{F}_{slow}$. When the query is easy, the frame does not need the switch adapter $\mathcal{S}_{adapter}$ and only output result $\mathcal{R}$ by $\mathcal{F}_{fast}$ as Equation 2.
@@ 2.2 Preliminary
Hierarchical Reasoning. We use a top-down scheme to reason over multi-scale granularity images effectively in order to reason and take advantage of world knowledge progressively. Similar to humans would look for some context clue to find specific objects relating to questions and zoom in if they think the answer lies in a particular region, we model this process with system switch adapters as Remark. 2.2 to focus on the context clue $\mathcal{C}_{clue}$ generated from the switch adapter as Equation 3.
We denote the MLLM as a proposal adapter $\mathcal{P}_{adapter}$ (visual agent). In System 2, FaST uses many visual agents to accomplish hierarchical reasoning. The frame tries to narrow down the search space by using the question $\mathcal{Q}$ and the previously obtained clue $\mathcal{C}_{clue}$ to let the proposal adapter output a region $Region$ that aligns with the question and the context clue as Equation 4.
@@
After getting the region, the visual agents $\mathcal{P}_{adapter}$ will be asked to focus on a more specific target with a bounding box $[Bboxes]$ complemented by the context clue $\mathcal{C}_{clue}$ and region $Region$ get from Equation 4. This process can reveal the step-by-step reasoning and be modeled as Equation 5.
@@
When we have a more specific target proposal(bounding box $[Bboxes]$), FaST will apply a fine-grained pixel-level mask decoder $\mathcal{P}_{Seg}$ as Equation 6 to output the specific mask part $[Mask]$ of the target proposal $[Bboxes]$ to focus on as Equation 6. We name this whole process from $Region$ to $[Mask]$ chain of evidence as Remark. 2.5 similar to thinking more and more deeply by humans.
@@
After getting the target proposal (bounding box $[Bboxes]$) from context clue $\mathcal{C}_{clue}$ with proposal adapter and specific mask part $[Mask]$ by missing objects with seg adapter, a chain of evidence is constructed as Remark 2.5 and Fig. 3. Our FaST framework then summarizes all this information ($\mathcal{I}$ and $\mathcal{Q}$) and the chain of evidence with switch adapter to give the final correct reasoning answer $Ans$ as Equation 7:
@@ 2.3 Implementation Details
$\bullet$ Chain of Evidence. When we apply the chain of evidence as Remark 2.5 in the LLM to get the answer as the final step like Equation 7. The whole sequence of the chain of evidence is too long to load in the $\mathcal{F}_{Slow}$. So FaST needs a visual sampler based on cross-attention that is trained to decrease the number of image tokens to a suitable length (from 256 to 32), apart from MLP projection.
@@ 4 Related works
LLM as Visual Agents. With the capabilities that LLMs have demonstrated in language understanding and generation (Ouyang et al., 2022; OpenAI, 2024; Zheng et al., 2023; Touvron et al., 2023a; b; Wang et al., 2024a; Hua et al., 2024; Mei et al., 2024; Jin et al., 2024a; Chen et al., 2024c), the research community has progressed to explore how LLMs can be enhanced with vision input for multimodal tasks as visual agents (Alayrac et al., 2022; Driess et al., 2023; Li et al., 2023a; Ge et al., 2023; Dai et al., 2024; Liu et al., 2023b; 2024a; Lin et al., 2024; Wang et al., 2024b). There are two paradigms for LLM-based visual agents: end-to-end based and tool-using visual agents. Following the principle of instruction tuning, end-to-end visual agents are trained with a curated visual instruction tuning dataset to digest features from multi-modality, unlocking the capability to answer visual questions (Huang et al., 2023; Luo et al., 2023; Zhu et al., 2023a; Bai et al., 2023; Zhang et al., 2023b; c; Chen et al., 2024a; Ye et al., 2023; Singh et al., 2019). For other visual tasks ($e.g.$, Segmentation, Detection, etc), end-to-end trained tailored agents can further perform downstream tasks (Pi et al., 2023; Peng et al., 2024; Lai et al., 2024; Chen et al., 2023b; Wang et al., 2023c; Dai et al., 2024; Wang et al., 2024c; 2023b; Jiang et al., 2023; Chen et al., 2023a; Zeng et al., 2024). Recent research has focused on leveraging improved vision encoders and fostering more detailed visual understanding, yielding promising results (Fan et al., 2024; Xu et al., 2024a; Shi et al., 2024). While these approaches can be implemented with direct instruction tuning data, they represent a ‘System 1’ type of training. This type of training primarily relies on the dataset’s quality and tends to provide direct answers that are prone to hallucinations, a consequence inherent to the nature of System 1 instruction tuning data. For the second paradigm, tool-using models are built on top of a frozen LLM with access to pretrained visual perception tools (Surís et al., 2023; Shen et al., 2023; Lu et al., 2023a). In this scenario, the LLM first selects visual tools and then decides by thoroughly analyzing the fine-grained information extracted by visual tools (Lu et al., 2023a; You et al., 2023; Wu et al., 2024). While external visual tools enhance the interpretability of the reasoning process, their complexity can introduce inaccuracies. Moreover, the abundance of information generated during reasoning may overshadow key details relevant to the query, resulting in incorrect answers.
@@
System 2 in AI. Recently, LLMs have been engineered to produce text that mimics the step-by-step reasoning process characteristic of human cognition, akin to the analytical and deliberate thought processes associated with what is termed as System 2 in the human cognition process (Qiao et al., 2023; Huang & Chang, 2023a; Wang et al., 2023a; Shaikh et al., 2023; Shao et al., 2024). The systematic approach to problem-solving is a hallmark across various domains, including mathematical word problems (Kojima et al., 2022; Wang et al., 2023d; Lightman et al., 2023; Cobbe et al., 2021; Liu et al., 2023c; Zhu et al., 2023b; Lu et al., 2023b), logical reasoning (Yao et al., 2023d; a; Besta et al., 2024; Wen et al., 2023; Lei et al., 2023; Cheng et al., 2024; Jin et al., 2024b), and multi-modal reasoning (Chen et al., 2024e; You et al., 2023; Wu & Xie, 2024). In Explainable AI, this systematic method is emulated by the model as it generates a text-based elucidation of its reasoning and decision-making process through step by step reasoning process ($e.g.$, chain of thought) (Han et al., 2024; Zhao et al., 2024; Jacovi & Goldberg, 2020; Hua & Zhang, 2022). However, it is crucial to recognize that LLMs, while powerful, are not exempt from encountering challenges when facing complex problems. One such challenge is the issue of hallucination (Zhou et al., 2024; Cui et al., 2023; Li et al., 2023b; Zhang et al., 2023a; Chen et al., 2024d; Guan et al., 2024a), which can distort the model’s reasoning process and lead to inaccuracies in the explanations provided. Initially, LLM reasoning is seen as only a linear chain of thoughts, where each step in the reasoning process is clearly articulated. As models evolve, they adopt more complex structures like hierarchical trees (Geng et al., 2023; Yao et al., 2023b) and intricate graphs (Besta et al., 2024), which enable them to handle much more complex problems but also restrict their general applicability because of increased topological complexity (Yao et al., 2023b; a; Besta et al., 2024; Lei et al., 2023; Wen et al., 2023). Moreover, these complex structures can lead to errors propagating through the model’s reasoning, causing a cascade of mistakes (Xu et al., 2024b). To counter this, incorporating feedback from intermediate reasoning steps and employing iterative refinement, which is similar to human reflection, could help mitigate errors (Chu et al., 2023; Tong et al., 2024b; Guan et al., 2024b; Madaan et al., 2023; Yuan et al., 2024; Wu & Xie, 2024). In unsupervised scenarios, such feedback is vital for enhancing the reasoning capabilities of LLMs and reducing errors (Yao et al., 2023c).
@@ 5 Conclusion
6 Acknowledgments
This research was supported by the National Science Foundation under Grant No.2242243. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of U.S. Naval Research Laboratory (NRL) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.